OpenStudy (anonymous):

Which pair of ratios forms a proportion? A:11 over 36, 121 over 180 B:13 over 21, 39 over 63 C:14 over 23, 70 over 92 D:20 over 51, 4 over 5

OpenStudy (anonymous):

B. \[\frac{13}{21}=\frac{39}{63}\to \frac{13}{21}=\frac{13}{21}\]